What is the current through the 62 resistor? 4V 22 w 3 V 62 himni 6 V
Answer #1

ingata! 3v. Find out- current through the bar resistore Solution - It Iz 145 A B 2 hIst Iz Lutz 3v IC Hem D 1 2 E IZ 6vLuole ABC DA - 4 + 6.1 3 +2[I, 712] = 0 - 4+ 61 -3 +2I, + 21 2 = 0 & I, + 212-720 8. I, + 2 1 2 = 7 -0 Lool DC EFD +3 - 61 -